OpenStudy (anonymous):

PLEASE HELP ME WILL I MEDAL AND FAN!!!!!!!!!!!!!!! The Andersons traveled 90 miles in 75 minutes. At this rate, how far would they travel in 2 hours? A. 144 mi B. 165 mi C. 180 mi D. 5400 mi

OpenStudy (anonymous):

@chosenmatt

OpenStudy (chosenmatt):

one sec

OpenStudy (anonymous):

is the ansewr 156.8

OpenStudy (lizz123):

90miles/75min is the same as 6miles/5min 2hours=120min, so if 120/5=24 you multiply 6miles*24=144 miles

OpenStudy (lizz123):

So your answer is A. 144

OpenStudy (anonymous):

thank you lizz123

OpenStudy (lizz123):

no problem @12hawks12

OpenStudy (anonymous):

can you @lizz123 help me in another one or @chosenmatt

OpenStudy (lizz123):

yeah wht is it

OpenStudy (anonymous):

Malorie can work 10 math problems in 3 minutes. At this rate, how many problems can she work in one hour? A. 30 B. 60 C. 120 D. 200

OpenStudy (lizz123):

\[\frac{ 10 }{ 30 \min } = \frac{ x }{ 60 \min }\]

OpenStudy (lizz123):

3x = 600 x= 200

OpenStudy (lizz123):

200 is your answer

OpenStudy (lizz123):

@12hawks12

OpenStudy (anonymous):

thanks

OpenStudy (lizz123):

no problem

OpenStudy (anonymous):

I have one more question

OpenStudy (lizz123):

yeah what

OpenStudy (anonymous):

Javi figures that 2 small pizzas will feed 3 people. He expects 15 people at the team pizza party. How many small pizzas will he need for 15 people? A. 6 B. 10 C. 30 D. 45

OpenStudy (lizz123):

is the question